A merchant blends tea that sells for $5.00 a pound with tea that sells for $2.00 a pound to produce 120 lb of a mixture that sells for $3.90 a pound. How many pounds of each type of tea does the merchant use in the blend?
Answer by richwmiller(17219)

a+b=120,
5*a+2*b=3.9*120
a=120-b
5*(120-b)+2*b=468
600-5b+2*b=468
-3*b=-132
b=44
a=120-b
a=76 lb at $5 b=44 lb at $2
check
5*76+2*44=3.9*120
380+88=468
468=468
ok